OpenStudy (hlambach):

Does not (17+r)^2 equal 289 +r^2?

OpenStudy (hlambach):

So the problem is r^2 + 24^2 = (17 + r)^2

OpenStudy (hlambach):

My teacher says it's is wrong.

OpenStudy (anonymous):

No, \[(a+b)² = a² + 2ab + b²\]. So, \[(17 + r)² = 17² + 2*17*r + r² = 279 + 34r + r²\]

OpenStudy (anonymous):

It's not 279 in my answer, its 289, sorry...

OpenStudy (hlambach):

whoops yeah i typed it wrong

OpenStudy (hlambach):

but wouldn't (r)^2 equal r^2?

OpenStudy (anonymous):

It depends. For example: \[(-3)² = 9\]\[-3² = -9\]

OpenStudy (hlambach):

where did the 2 x 17 x r come from?

OpenStudy (hlambach):

oh

OpenStudy (anonymous):

Like i said, \[(a+b)² = a² + 2ab + b²\] Changing a to 17 and b to r, you wil have:\[(17 + r)² = 17² + 2*17*r + r²\]

OpenStudy (hlambach):

oh... i guess that makes sense...

OpenStudy (anonymous):

You can see doing: \[(17 +r)²=(17+r)(17+r)=17*17+17r+17r+r² = 289 + 34r +r²\]

OpenStudy (hlambach):

ah! got it now! thanks!